# Fast and Globally Consistent Normal Orientation based on the Winding Number Normal Consistency

### v1

@@ 5.2.7 Real-world Scans
Fig. 13 shows the reconstruction results using normals estimated by our method, iPSR [19] and Dipole [34]. To adapt to scanner noise, we set $w_{1}=0.01$ and $w_{2}=0.04$ for our method, and set the point weight to 0.5 for iPSR and the final reconstruction step. Dipole still suffers from severe misorientation (see the Armadillo head and the Dragon tail). Even though iPSR can infer a mostly correct orientation, it is less capable of handling scanner noise. Our method not only produces globally consistent orientations but also achieve a good balance between noise handling and detail preserving.
5.3 Complexity Analysis
@@
In this section we study one important aspect of our method: the performance improvement in terms of space/time complexity. Considering the complexity of the baselines, we divide the experiments into three groups: small-scale (1,000$\sim$5,000 points), medium-scale (10,000$\sim$50,000 points) and large-scale (100,000$\sim$1,000,000 points). For the small-scale study, we experiment with all methods. We exclude PGR from large-scale experiments, and exclude GCNO from both medium-scale and large-scale experiments. For each experiment, we use a Armadillo model uniformly sampled without noise.
@@
Furthermore, to present a more comprehensive analysis, we break down each method into a preprocessing stage and a main stage. For our method, iPSR and PGR, the preprocessing stage includes octree building, and the main stage includes normal solving. For GCNO, preprocessing includes Delaunay triangulation and the main stage includes the LBFG-S iterations. For Dipole, preprocessing incldues all steps prior to inter-patch propagation, and the main stage includes everything else. Note that the final reconstruction stage and file I/O are excluded for all methods.
@@
In Fig. 14, we show the total running time $T_{\rm total}=T_{\rm pre}+T_{\rm main}$, preprocessing time $T_{\rm pre}$, main stage running time $T_{\rm main}$ and the total peak memory (CPU plus GPU). Time and memory measurements are shown in log-scale due to large variations between different methods. In terms of total running time, it can be seen that our method is the fastest for medium and large inputs, and even our CPU implementation (with OpenMP) is faster than other methods. Note that for small-scale inputs the CPU implementation is faster due to the overhead associated to CUDA runtime library and cross-device data transfer. In terms of memory consumption, our method is slightly larger than iPSR and GCNO for small and medium inputs, mainly because our implementation is integrated with PyTorch [2, 35] as an extension module, which involves PyTorch and CUDA overheads. However, as the input size scales up, our method has the lowest memory footprint.
@@
Additionally, we would like to highlight our time and memory consumption improvements over the other two winding number-based methods: PGR and GCNO. Both of PGR and GCNO requires repeated evaluations of winding numbers, and both do so in a naive way by directly summing all contributions from each input point, leading to an $N^{2}$ complexity lower-bound for each evaluation ($N$ being the number of points). More specifically, to orient a point cloud with 100,000 points, PGR requires at least 3000G momery and GCNO would takes weeks. On the other hand, our treecode-based algorithm is known to have $O(N\log N)$ complexity [4]. Thus, our method easily scales up to millions of points and can finish normal orientation for these large-scale inputs in minutes.
@@
Figure 14: Complexity comparisons of different methods, evaluated on a uniformly sampled Armadillo model with different sample numbers (first row: small-scale; second row: medium-scale, third row: large-scale). We show the total running time $T_{\rm pre}+T_{\rm main}$, preprocessing time $T_{\rm pre}$, main-stage execution time $T_{\rm main}$ and peak memory usage (MB).
5.4 Evaluation of Individual Components
In this section, we evaluate individual components of our algorithm: the smoothing width in Sec. 4.4 and the iterative steps in 4.3.
5.4.1 Smoothing width
As mentioned in previous works [32, 31] that also apply smoothing modifications to the winding number formula, this modification allows ignoring local points, leading to better robustness to noise or non-uniformity. Intuitively, a large smoothing width produces more consistent orientation but may result in oversmooth, and a small smoothing width preserves more details but may fail to orient normals consistently. In contrast to using an invariant smoothing width as in prior work [31], our width scheduling scheme that starts at a large width and ends with a small width achieves both consistent orientation and detail preserving. The first row in Fig. 15 compares using invariant smoothing widths and decreasing smoothing widths. As is evident from Fig. 15, a small width leads to orientation errors in some concave regions while a large width causes oversmoothing. Our scheduling scheme achieve consistent and detailed normal estimation.
@@ 5.4.1 Smoothing width
Figure 15: Study of the effect of the smoothing width. First row: decreasing the width during the iterative process allows both consistent orientation and detail preserving. Second row: adjusting widths allows explicit control over the smoothness of normals.
5.4.2 Steps in the Iterative Algorithm
@@
Figure 16: Ablation study of the effectiveness of WNNC updates. We show angular error maps of normals estimated with (w/) and without (w/o) using the WNNC updates. The later introduces notably larger angular error especially in concave regions.
6 Discussion
6.1 Conclusion
@@
In this work, we propose a novel property, namely the winding number normal consistency (WNNC), derived from the winding number formula to address the problem of estimating globally consistent normal orientations. We turn the WNNC property into a practical, efficient yet embarrassingly simple iterative algorithm. We also observe the special structures in the iterative algorithm that allows acceleration using the treecode algorithm [4]. Lacking an existing package of the treecode algorithm that suits our needs (GPU-based and supporting our operators), we implement our own CUDA kernels. With extensive experiments on a variety of datasets, we exhibit the superior performance of our algorithm and implementation over recent state-of-the-art methods, in terms of both normal orientation quality and computational costs.
6.2 Limitations and Future Work
@@
The normal estimation quality of our method still highly depends on the smoothing width. As explained in Sec. 5.4.1, an improper choice can lead to either wrong orientation or ovesmoothing. We empirically found that the widths can be kept the same for data with similar noise levels and structural complexity. Hence, adjusting the width parameters shouldn’t take much effort. However, it is still desirable for future work to explore how to automatically determine the optimal width parameters.
@@
Currently we use the treecode algorithm to accelerate winding number related evaluations because it is easy to implement and parallelize. A possible future direction is to implement a more advanced acceleration technique, the fast mulipole method (FMM) [14], which further reduces complexity and allows precision control. However, FMM requires more careful mathematical analysis and is more difficult to implement and parallelize. Thus, we leave it as future work.
Figure 17: Our method is built on the winding number formula, which cannot handle non-manifold surfaces.

### v2

@@ 5.2.6. Real-world Scans
Fig. 12 shows the reconstruction results using normals estimated by our method, iPSR (Hou et al., 2022) and Dipole (Metzer et al., 2021). To adapt to scanner noise, we set $w_{1}=0.01$ and $w_{2}=0.04$ for our method, and set the point weight to 0.5 for iPSR and the final reconstruction step. Dipole still suffers from severe misorientation (see the Armadillo head and the Dragon tail). Even though iPSR can infer a mostly correct orientation, it is less capable of handling scanner noise. Our method not only produces globally consistent orientations but also achieve a good balance between noise handling and detail preserving.
5.3. Convergence Analysis
In this section we compare the convergence processes of our method and the baselines. In Fig. 13 and Fig. 14, we present the qualitative and the quantitative results for the intermediate convergence steps using two models: 3DBenchy with 8,000 points and XYZRGB Statuette with 50,000 points. For fairness, we show the first 40 iterations for all iterative methods. Note that our method is able to achieve an overall correct orientation in the first few iterations. Quantitatively, our method also achieves lower normal error than other baselines after convergence.
5.4. Complexity Analysis
@@
In this section we study one important aspect of our method: the performance improvement in terms of space/time complexity. Considering the complexity of the baselines, we divide the experiments into three groups: small-scale (1,000$\sim$5,000 points), medium-scale (10,000$\sim$50,000 points) and large-scale (100,000$\sim$1,000,000 points). For the small-scale study, we experiment with all methods. We exclude PGR from large-scale experiments, and exclude GCNO from both medium-scale and large-scale experiments. For each experiment, we use an Armadillo model uniformly sampled without noise.
@@
Furthermore, to present a more comprehensive analysis, we break down each method into a preprocessing stage and a main stage. For our method, iPSR and PGR, the preprocessing stage includes octree building, and the main stage includes normal solving. For GCNO, preprocessing includes Delaunay triangulation and the main stage includes the LBFG-S iterations. For Dipole, preprocessing includes all steps prior to inter-patch propagation, and the main stage includes everything else. Note that the final reconstruction stage and file I/O are excluded for all methods.
@@
In Fig. 15, we show the total running time $T_{\rm total}=T_{\rm pre}+T_{\rm main}$, preprocessing time $T_{\rm pre}$, main stage running time $T_{\rm main}$ and the total peak memory (CPU plus GPU). Time and memory measurements are shown in log-scale due to large variations between different methods. In terms of total running time, it can be seen that our method is the fastest for medium and large inputs, and even our CPU implementation with OpenMP is faster than other methods. Note that for small-scale inputs the CPU implementation is faster due to the overhead associated to CUDA runtime library and cross-device data transfer. In terms of memory consumption, our method is slightly larger than iPSR and GCNO for small and medium inputs, mainly because our implementation involves PyTorch and CUDA overheads. However, as the input size scales up, our method has the lowest memory footprint.
@@
Additionally, we would like to highlight our time and memory consumption improvements over the other two winding number-based methods: PGR and GCNO. Both of PGR and GCNO requires repeated evaluations of winding numbers, and both do so in a naive way by directly summing all contributions from each input point, leading to an $N^{2}$ complexity lower-bound for each evaluation ($N$ being the number of points). More specifically, to orient a point cloud with 100,000 points, PGR requires at least 3000 GB memory and GCNO would take weeks. On the other hand, our treecode-based algorithm is known to have $O(N\log N)$ complexity (Barnes and Hut, 1986). Thus, our method easily scales up to millions of points and can finish normal orientation for these large-scale inputs in minutes.
@@
Figure 15. Complexity comparisons of different methods, evaluated on a uniformly sampled Armadillo model with different sample numbers (first row: small-scale; second row: medium-scale, third row: large-scale). We show the total running time $T_{\rm pre}+T_{\rm main}$, preprocessing time $T_{\rm pre}$, main-stage execution time $T_{\rm main}$ and peak memory usage (MB).
[…]
Complexity analysis of different methods.
5.5. Evaluation of Individual Components
In this section, we evaluate individual components of our algorithm: the smoothing width in Sec. 4.4 and the iterative steps in Sec. 4.3.
5.5.1. Smoothing width
As mentioned in previous works (Lu et al., 2018; Lin et al., 2022) that also apply smoothing modifications to the winding number formula, this modification allows ignoring local points, leading to better robustness to noise or non-uniformity. Intuitively, a large smoothing width produces a more consistent orientation but may result in oversmoothing, and a small smoothing width preserves more details but may fail to orient normals consistently. In contrast to using an invariant smoothing width as in prior work (Lin et al., 2022), our width scheduling scheme that starts at a large width and ends with a small width achieves both consistent orientation and detail preserving. The first row in Fig. 16 compares using invariant smoothing widths and decreasing smoothing widths. As is evident from Fig. 16, a small width leads to orientation errors in some concave regions while a large width causes oversmoothing. Our scheduling scheme achieve consistent and detailed normal estimation.
@@ 5.5.1. Smoothing width
Figure 16. Study of the effect of the smoothing width. First row: decreasing the width during the iterative process allows both consistent orientation and detail preserving. Second row: adjusting widths allows explicit control over the smoothness of normals.
[…]
We remark that we also tried other scheduling schemes (e.g., cosine, exponential) but found no obvious difference, as long as the width drops from a large $w_{2}$ to a small $w_{1}$ smoothly. Thus, we chose the most simple linear scheduling. Furthermore, due to the width scheduling scheme, we do not adopt any early termination strategy because it may stop at a large $w$, resulting in oversmoothed normals. Note that using a fixed number of iterations has little practical impact, because running full 40 iterations with our method is still faster than 1 iteration of the baselines.
To provide a more intuitive view of the smoothing width, we report the average number of neighbors in the $w$-neighborhood of a 100,000-point uniformly sampled Armadillo model in Fig. 17.
$w$$N_{\rm nbr}$
Figure 17. Plot of the average number $N_{\rm nbr}$ of neighbor points in the $w$-neighborhood of each point (including itself) in a uniformly sampled Armadillo model with 100,000 points. The average point spacing is $0.0076$. All the above statistics are computed after normalizing to the unit cube.
[…]
Plot of the average number of neighbors points in a $w$-neighborhood.
5.5.2. Steps in the Iterative Algorithm
@@
Figure 18. Ablation study of the effectiveness of WNNC updates. We show angular error maps of normals estimated with (w/) and without (w/o) using the WNNC updates. The latter introduces notably larger angular error especially in concave regions.
[…]
Ablation study of the WNNC updates.
6. Discussion
6.1. Extended Discussions
In this section we conduct extended discussions to provide more insights into our method.
6.1.1. Why not use estimated local areas during optimization?
Using estimated local surface element areas $\sigma_{i}$ adds constraints $|\mu_{i}|=\sigma_{i}$ to Eq. (15), making the feasible space of $\mu$ non-linear. Consequently, Algorithm 2 must be replaced by a non-linear solver, which is not immediately or easily possible using our operators. As an alternative, assuming known/estimated local areas $\sigma_{i}$, we can enforce $|\mu_{i}|=\sigma_{i}$ by changing the rescaling step in Algorithm 3 to
(22) $\mu_{i}\leftarrow\hat{\mu}_{i}(\sigma_{i}/|\hat{\mu}_{i}|).$
In Fig. 19, we use the GT mesh to compute Voronoi areas, and proceed with Eq. (22). As Fig. 19 shows, using fixed local areas leads to notable orientation errors. This suggests even with accurate local areas, our algorithm cannot benefit from using fixed areas because the feasible space becomes non-linear. Furthermore, in practice, local area estimation can be difficult or inaccurate especially for noisy inputs or sharp features. In summary, using fixed estimated areas makes the solution space non-linear, and makes Eq. (15) harder to solve, is less robust to noise, leads to extra computational complexity and brings no obvious benefit. Hence, we do not use estimated local areas.
GT normalsOriginal methodUsing estimated $\sigma_{i}$and Eq. (22)
Figure 19. We test the alternative where each local surface element is enforced to have a pre-estimated area. Experiments show this will lead to notable orientation errors (the cave-like part in the middle). Points rendered here are colored by their normals.
[…]
Experiments of whether our method benefits from estimated surface areas.
6.1.2. Do solved local areas converge to actual local areas?
Not necessarily. In Table 5, we conduct an experiment on the area convergence using a unit sphere model. We take its vertices as the input point cloud (163,842 points), whose local Voronoi areas $\{\sigma_{i}^{\rm GT}\}$ are directly computed from the GT triangulation. We empirically find after 40 iterations when the normals are already accurate, solved local areas $\{|\mu_{i}|\}$ do not accurately match GT areas $\{\sigma_{i}^{\rm GT}\}$, while the $k$NN-based area estimation method (Barill et al., 2018) is more accurate, as shown in Table 5. We remark that this work does not claim area convergence, and whether areas converge or not is irrelevant, because area convergence is not required for accurate normal convergence.
Table 5. A comparison of the statistics for GT surface element Voronoi areas $\{\sigma_{i}^{\rm GT}\}$, estimated areas $\{\sigma_{i}^{k\rm NN}\}$ using $k$ nearest neighbors as in Barill et al. (2018), and solved areas $\{|\mu_{i}|\}$ for a uniformly sampled unit sphere model.
$\{\sigma_{i}^{\rm GT}\}$$\{\sigma_{i}^{k\rm NN}\}$$\{|\mu_{i}|\}$
Min $6.795\times 10^{-5}$ $6.794\times 10^{-5}$ $7.465\times 10^{-5}$
Max $9.253\times 10^{-5}$ $9.252\times 10^{-5}$ $8.330\times 10^{-5}$
Mean $7.670\times 10^{-5}$ $7.670\times 10^{-5}$ $7.858\times 10^{-5}$
Total $12.566$ $12.566$ $12.875$
Initial normalsw/o grad. stepw/ grad. stepZero init.One side flippedRandom init.GT normals
Figure 20. Normal orientation results obtained by different normal initialization strategies and whether the gradient step (Algorithm 2) is used.
[…]
Normal orientation results obtained by different normal initialization strategies and whether the gradient step is used.
GT meshOursPGR normals+SPSRPGR normals+WNF
Figure 21. Comparisons between our method and PGR using different reconstruction methods. Our method outperforms PGR regardless of whether it uses Screened Poisson Surface Reconstruction (SPSR) or the winding number field-based reconstruction (WNF).
[…]
Comparisons with PGR in a more fair setting.
6.1.3. Can WNNC alone converge to consistent normals?
No. Note that a consistently outward normal vector field and a consistently inward vector field are both invariant up to scaling under WNNC updates, which suggests WNNC updates alone cannot handle the inside/outside ambiguity. Furthermore, we empirically found two key factors in the convergence of our algorithm: zero initialization and the gradient steps w.r.t. Eq. (15). In Fig. 20, we experiment with different initialization strategies and whether Eq. (15) is applied. As Fig. 20 shows, if the normals are not zero-initialized, the normals cannot converge to a consistent orientation regardless of whether Eq. (15) is used. On other hand, if we only apply WNNC updates to zero-initialized normals without Eq. (15), the optimization would simply stay at zero vectors according to Algorithm 3. Good convergence is only possible with both zero initialization and Eq. (15).
6.1.4. WNNC vs. PGR
While the effectiveness of our method seems to heavily rely on the objective Eq. (15) proposed by PGR (Lin et al., 2022), the latter is fundamentally built on an underdetermined system (Eq. (15)). As discussed in Lin et al. (2022), even if given enough iterations, their solved normals usually do not align well with GT normals, resulting in low normal accuracy. On the other hand, our WNNC formulation provides a new set of constraints for normal directions, leading to a more determined system. By alternating gradient steps w.r.t. Eq. (15) and WNNC updates, our method quickly converges to an accurate and consistently oriented normal vector field. As shown in Sec. 5.3, our method achieves much higher accuracy given the same number of iterations. Other comparisons also show our method achieves better accuracy after full convergence.
Furthermore, the authors of PGR (Lin et al., 2022) suggested that the normals solved by PGR are not really compatible with SPSR (Kazhdan and Hoppe, 2013), and better reconstructions can be obtained by plugging these normals back into the winding number field (WNF). While we have previously used SPSR as the reconstruction for all methods for fairness, we also test PGR with WNF-based reconstruction as in Lin et al. (2022). In Fig. 21, we observe that our method still outperforms PGR despite the final reconstruction method.
6.2. Conclusion
@@
In this work, we propose a novel property, namely the winding number normal consistency (WNNC), derived from the winding number formula to address the problem of estimating globally consistent normal orientations. We turn the WNNC property into a practical, efficient yet embarrassingly simple iterative algorithm. We also observe the special structures in the iterative algorithm that allows acceleration using the treecode algorithm (Barnes and Hut, 1986). Lacking an existing package of the treecode algorithm that suits our needs (GPU-based and supporting our operators), we implement our own CUDA kernels. With extensive experiments on a variety of datasets, we exhibit the superior performance of our algorithm and implementation over recent state-of-the-art methods, in terms of both normal orientation quality and computational costs. Furthermore, our implementation is integrated with PyTorch (Ansel et al., 2024; Paszke et al., 2019) and is easy to use. We hope our implementation can facilitate future research into winding numbers.
Figure 22. Our method is built on the winding number formula, which cannot handle non-manifold surfaces.
[…]
Failure case of open, non-manifold surfaces.
6.3. Limitations and Future Work
@@
The normal estimation quality of our method still highly depends on the smoothing width. As explained in Sec. 5.5.1, an improper choice can lead to either wrong orientation or oversmoothing. We empirically found that the widths can be kept the same for data with similar noise levels and structural complexity. Hence, adjusting the width parameters shouldn’t take much effort. However, it is still desirable for future work to explore how to automatically determine the optimal width parameters.
@@
Currently, we use the treecode algorithm to accelerate winding number related evaluations because it is easy to implement and parallelize. A possible future direction is to implement a more advanced acceleration technique, the fast multipole method (FMM) (Greengard and Rokhlin, 1997), which further reduces complexity and allows precision control. However, FMM requires more careful mathematical analysis and is more difficult to implement and parallelize. Thus, we leave it as future work.
Accelerating existing winding number-based algorithms
Another line of future work would be to extend our treecode implementation to accelerate other existing methods that use the winding number field, e.g. GCNO (Xu et al., 2023). Furthermore, note that we have implemented the $A^{T}$ operator, which is exactly what is needed for differentiating through the winding number formula. This means our implementation can also be incorporated into differentiable programming involving winding numbers. Since our implementation has been incorporated with the PyTorch (Ansel et al., 2024; Paszke et al., 2019) framework, differentiable programming with winding numbers can be made quite easy.
A theoretical proof of convergence
While our proposed method exhibits notable superiority over other baselines, our solution remains an engineering solution without a rigorous theoretical proof of convergence. Future work should explore the theoretical aspects of the WNNC formulation.
Acknowledgements.
The work is supported by the National Science Foundation of China under Grant Number Grant #62125107 and Grant #92370125.